OpenStudy (anonymous):

trying to simplify this expression (x^2y^6)1/2 over (-2xy)^3

OpenStudy (anonymous):

-( (y^3) / (16x) )

OpenStudy (anonymous):

Not understanding how you got this answer.

OpenStudy (anonymous):

\[ (x^2 y^6)^\frac12 \] simplifies to \[ x y^3 \]

OpenStudy (anonymous):

\[ (-2xy)^3 = -8x^3y^3 \]

OpenStudy (anonymous):

and \[ \frac{xy^3}{-8x^3y^3} = \frac{1}{-8x^2} \]

OpenStudy (anonymous):

so that would be your answer